### v1

@@ 3 Arguments
As a result, there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels. Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multiscale causal learning. We aren’t suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
Dropping Hume’s Guillotine: Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just abstractions we ascribe to the behaviour of the lowest level collective). Previous work argues learned behaviour must be motivated by attraction and repulsion from aspects of the environment [bennett2023c]. This eventually culminates in “feelings” at higher levels [bennett2023d]. Assuming that is true, then by intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that impetus. We might describe this as “dropping Hume’s Guillotine” to separate is from ought. We suggest that a computer is not a “syntactic engine”, as it is sometimes portrayed [kim2011]. Syntax has meaning to the computer, otherwise it would do nothing. However the “meaning” of an instruction is what it does, just as the “meaning” of a declarative program is its truth conditions. What a computer is, is meanings realised in circuitry by human hands. That is why it “interprets” syntax as it does. We speculate one reason computers may not “feel” and “experience” meaning as a human would is because there is a static, hard coded abstraction layer that allows high level goal directed behaviours to exist in the absence of the low level goal directed behaviours of which they are an emergent property. As a result, interpretation is fixed rather than fluid and motivated (e.g. by affect).
[…]

### v2

@@ 3 Arguments and Results
As a result, there is no adaptive system underpinning the ISA which can adjust policy at lower levels to facilitate goal directed behaviour at higher levels. Intuitively, if an abstraction layer is a window onto the world, then utility measures how “good” that view is for a task. A static abstraction layer is a window that cannot move, while an MCA can change policy at lower levels, to move the window into an ideal location. When we create an ISA we insert a static abstraction layer at an intermediate level, so the vocabulary at the highest levels is limited to only compositions and permutations of those abstractions we have embodied in the ISA. In comparison, an MCA can learn policies at lower levels, to change the vocabulary of higher levels and facilitate multiscale causal learning. We aren’t suggesting lower levels only optimise for utility at higher levels. After all, an increase in utility may increase the computational cost of searching for correct policies. Furthermore, a spatially extended environment constrains the size of vocabularies at all levels [16], because bounded systems can contain only finite information [32]. We mean lower levels learn an inductive bias for higher levels, balancing utility with other factors in service of goals at every scale.
3.1 Precondition for Collective Identity
According to Cao and Yamins [14] tasks are more “difficult” when they have fewer correct policies, and “easier” when they have many correct policies. In our formalism, Cao and Yamins’ position implies $\mathfrak{v}$-task $\alpha$ is “easier” than $\omega$ if $\mid\Pi_{\alpha}\mid>\mid\Pi_{\omega}\mid$. Collectives that exhibit high level goal directed behaviour that is “easier” have more correct high level policies to choose from. If the policy at a lower level $i-1$ changes such that the task at the higher level $i$ becomes “harder”141414i.e. by reducing $\mid\Pi^{i}\mid$ at the higher level $i$., then this could result in their being no correct policy at level $i$. If there are no longer any correct policies at level $i$ after a change in policy at level $i-1$, and if the system is robust and exhibits graceful degradation, then part of a collective may “break off” (loosening the constraint on correct policies, for both the breakaway part and the part left behind). In more formal terms, say we have a collective of cells $\lambda^{i}(\mathfrak{v}^{i})$ that includes a cell $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i})$ s.t. $\mathfrak{f}(E_{\pi^{i-1}_{\mathfrak{a}}})=\mathfrak{v}^{i}_{\mathfrak{a}}$. Now assume the policy $\pi^{i-1}_{\mathfrak{a}}$ changes to $\pi^{i-1}_{\mathfrak{b}}$ s.t. $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$. A subset of the collective may continue functioning as before with a shared policy, but doing so requires discarding a part of the collective which cannot adopt that policy. This is akin to “selective forgetting” [11] to cope with false negatives, but instead of forgetting data inconsistent with an otherwise good policy, we would “discard” a child task $\alpha\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ of $\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ when $\Pi_{\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})}=\emptyset$, so that we obtain a task $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})\sqsubset\lambda^{i}(\mathfrak{v}^{i}_{\mathfrak{b}})$ s.t. $\alpha\not\sqsubset\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ and $\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})$ has correct policies, meaning $\Pi_{\lambda^{i^{\prime}}(\mathfrak{v}^{i}_{\mathfrak{b}})}\neq\emptyset$. $\alpha$ is now isolated from the informational structure of the collective and so may have an extremely broad set of “correct” policies to choose from.
This is analogous to cancer as cells isolated from a collective informational structure, that lose their high level identity, and pursue lower level goals [27, 28]. If $\alpha$ is a cell, it can now revert to primitive transcriptional behaviours and act like cancer.
4 Discussion
We have formalised MCAs and shown the capacity for WPO at high levels implies WPO at low levels. Causal identities are constructed at different scales. Hence we call it multiscale-causal-learning (MCL). This seems to indicate MCA-like designs could result in more efficient and reliable AI in future.
4.1 Dropping Hume’s Guillotine
Given how we have formalised MCAs, the lowest level is all higher levels (or to put it another way, all higher levels are just classifications we ascribe to parts of the behaviour of the whole). Previous work argues learned behaviour must inevitably be motivated by attraction and repulsion from aspects of the environment [11]. By intervening in the causal chain to create a static abstraction layer, we disconnect higher levels from that motivating impetus. We might describe this as “dropping Hume’s Guillotine” to separate is from ought. This relates to the hard problem of consciousness [34]. According to our formalism, all abstractions constructed by living things are causal identities “goal directed” in the sense of being constructed in response to motivating affect. In dropping Hume’s Guilotine, high level human abstractions would become disconnected from the MCA that constructed them. It may be that, to a human or other MCA, all information amounts to statements of “ought” (motivated, value judgements) [15]. Purely representational content (free of any goal directed motive or connotation) may exist only in the static abstraction layers we engineer.
4.2 Future Research
4.2.1 AI-Safety, Distributed and Cyberphysical-Systems
The UK’s Advance Research and Invention Agency recently initiated a program called “Safeguarded AI” [35]; an ambitious research project that aims to erect a cyberphysical system meeting the standards laid out in Guaranteed Safe AI (GS) [36]. This involves constructing advanced general AI of level ASL-4 [37] [^15] Meaning an artificial general intelligence [38] capable of autonomous research and development that can evaluate the outcomes of deploying less advanced AI, and safeguard humanity against “rogue” AI. in a cyberphysical system. That system seems intended to safeguard against threats from both within and without. The idea of a “rogue AI” from within this system is analogues to cancer, adopting an “unsafe” policy outside the scope of what is permitted by collective. This suggests we could formalise such a system as an MCA, in order to prove properties relevant to AI safety. Control of cyberphysical and distributed systems in general might be framed in terms of an MCA, in order to better understand how different levels in a hierarchy or different parts of a horizontal system constrain each-other.
4.2.2 Active Inference
The sample efficiency of organoids [8] reveals what active inference misses; WPO. Future research might integrate WPO with active inference. Surprise minimisation [7] could be considered an extremely weak policy in and of itself. However, what really matters is fitness [15], not surprise. In adopting machine learning as its foundation, active inference may have inadvertently dropped Hume’s Guillotine, emphasising the wrong goal.